A swimmer wishes to cross a 500m river flowing at 5km/h. His speed with respect to water is 3km/h. The shortest possible time to cross the river is 5n min. Find the value of n?

Grade:12th pass

1 Answers

Ayush maurya
5 years ago
Given Vr=5km/h and Vbr=3km/h
  1. t=w/Vbr sin alpha
  2. Sinalpha=Vr/Vbr
Find the sinalpha=5/3 
Alpha=37° 
Put in 1 
5n=500/3×5/3
N=500